In recent years, Concord Transportation purchased three used buses. Because of frequent turnover in the accounting department, a different accountant selected the depreciation method for each bus, and various methods were selected. Information concerning the buses is shown as follows. Bus Acquired Cost Salvage Value Useful Life in Years Depreciation Method 1 1/1/18 $ 96,600 $ 7,000 4 Straight-line 2 1/1/18 123,000 12,000 5 Declining-balance 3 1/1/19 78,100 8,500 5 Units-of-activity For the declining-balance method, the company uses the double-declining rate. For the units-of-activity method, total miles are expected to be 116,000. Actual miles of use in the first 3 years were 2019, 26,000; 2020, 36,000; and 2021, 30,500.
Answer #1

For the first bus annual depreciation will be = (96600 - 7000)/4 = $22,400 per year

Year Beginning book value Depreciation Ending book value
2018                                    96,600                  22,400                                    74,200
2019                                    74,200                  22,400                                    51,800
2020                                    51,800                  22,400                                    29,400
2021                                    29,400                  22,400                                      7,000

2nd bus: Here rate as per straight line method = 100%/5 years = 20% per year. So double declining rate =20%*2 = 40% per year

Year Beginning book value Depreciation rate Depreciation Ending book value
2018                            123,000.00 40%                              49,200.00                       73,800.00
2019                              73,800.00 40%                              29,520.00                       44,280.00
2020                              44,280.00 40%                              17,712.00                       26,568.00
2021                              26,568.00 40%                              10,627.20                       15,940.80
2022                              15,940.80                                3,940.80                       12,000.00

3rd bus: Total miles = 116,000. Thus depreciation per mile = (78100-8500)/116000 = $0.60 per mile

A B C D = B*C
Year Beginning book value Mile Depreciation per mile Total depreciation Ending book value
2019                              78,100.00                     26,000.00                                        0.60                       15,600.00                       62,500.00
2020                              62,500.00                     36,000.00                                        0.60                       21,600.00                       40,900.00
2021                              40,900.00                     30,500.00                                        0.60                       18,300.00                       22,600.00